OpenStudy (anonymous):

what would be the increasing and decreasing intervals of the equation y= 1/3x(squared)+4x+12

OpenStudy (campbell_st):

is this calculus or algebra...?

OpenStudy (anonymous):

This is from my trigonometry class (algebra)

OpenStudy (campbell_st):

A simple method is to decide is the parabola concave up or concave down Concave up decreasing to the left of the vertex increasing to the right of the vertex Concave Down increasing to the left og the vertex decreasing to the right of the vertex. so you need to find the x value in the vertex... use the line of symmetry for a parabola \[y=ax^2 + bx + c\] the line of symmetry is\[x = \frac{-b}{2 \times a}\] you have b = 4 and a = 1/3 substitute them to find the line of symmetry

OpenStudy (anonymous):

OK thanks!